The length of a rectangle is 4 centimeters less than its width. What are the dimensions of the rectangle if its area is 192 square centimeters.
Answer by Cromlix(4381)

Hi there,
Make width = x
Length = 4 - x
Area = length x width
192 = (x - 4)x
192 = x^2 - 4x
x^2 - 4x - 192 = 0
(x + 12)(x - 16) = 0
x = 12 = 0
x = -12 (Disregard as -ve)
x - 16 = 0
x = 16
width = 16 centimeters
Length = 12 centimeters
Hope this helps :-)